OpenStudy (anonymous):

Please help: 2x-7y=52 6x+44=y solve by substitution method

OpenStudy (anonymous):

so u already have y so some of the work is done. Just plug it into the first equation.

OpenStudy (anonymous):

2x-7(6x+44)=52

OpenStudy (anonymous):

solve for x.

OpenStudy (anonymous):

I have a lot of trouble with linear equations. I don't understand how to do them. I need step by step instructions, but in college, I guess they figure you should already understand the basics, but I don't. :(

OpenStudy (anonymous):

2x-42x-308=52

OpenStudy (anonymous):

combine like terms.

OpenStudy (anonymous):

-40x-308=52

OpenStudy (anonymous):

-40=308+52

OpenStudy (anonymous):

-40x=360

OpenStudy (anonymous):

divide by -4

OpenStudy (anonymous):

x=360/-4

OpenStudy (anonymous):

x=-90

OpenStudy (anonymous):

we solved for x, now lets solve for y.

OpenStudy (anonymous):

y=6x+44

OpenStudy (anonymous):

plug in x: y=6(-90)+44

OpenStudy (anonymous):

y=-540+44

OpenStudy (anonymous):

good job mate

OpenStudy (anonymous):

y=-496

OpenStudy (anonymous):

thanks.

OpenStudy (anonymous):

so the answer should be (-90,-496)

OpenStudy (anonymous):

Thank you so, so much. I think I'm finally begining to grasp the concept. :)

OpenStudy (anonymous):

I may be back with more though if I run into a problem I just can't figure out. :)

OpenStudy (anonymous):

no problem. Good luck on your math studies.

OpenStudy (anonymous):

Thank you, again.